what is the value of sin 4530
 

Grade:10

2 Answers

srikanth
9 years ago
-1/2
Utkarsh Gupta
9 years ago
Here we get some rotations of 360 +330 extra left.
4530%360=330
 
the 220 left will be effective angle as argument....
 
sin 330=-sin(360-330)=-1/2;